OpenStudy (anonymous):

Let f(x)=5x and g(x)=25x^2/sqrt 5(x-2)

OpenStudy (anonymous):

Write g(x) in terms of f(x)

ganeshie8 (ganeshie8):

f(x) = 5x divide 5 both sides f(x)/5 = x which is same as x = f(x)/5

ganeshie8 (ganeshie8):

plug that in g(x) in place of "x" everywhere

OpenStudy (anonymous):

can you help me

ganeshie8 (ganeshie8):

write down g(x) first : \[g(\color{blue}{x})=\dfrac{25\color{blue}{x}^2}{\sqrt{5(\color{blue}{x}-2)}}\]

ganeshie8 (ganeshie8):

simply replace \(\color{blue}{x}\) by \(\color{blue}{f(x)/5}\) and simplify

ganeshie8 (ganeshie8):

\[\begin{align}g(\color{blue}{x})&=\dfrac{25\color{blue}{(f(x)/5)}^2}{\sqrt{5(\color{blue}{f(x)/5}-2)}}\\~\\ &=\dfrac{25\color{blue}{(f(x))^2/25}}{\sqrt{\color{blue}{f(x)}-10}}\\~\\ &=\dfrac{\color{blue}{(f(x))^2}}{\sqrt{\color{blue}{f(x)}-10}}\\~\\ \end{align}\]

OpenStudy (anonymous):

so that's the answer the last step right?

ganeshie8 (ganeshie8):

Yes :/

OpenStudy (anonymous):

thank you so much(; Can you help me with one more similar one

ganeshie8 (ganeshie8):

Id love to help, but if it is a similar one, I'm pretty sure you can work it on your own :)

ganeshie8 (ganeshie8):

give your best try

OpenStudy (anonymous):

Please.....Let f(x) =2sqrt x and g(x)=x^2+x. Write g(x) in terms of f(x)

ganeshie8 (ganeshie8):

take first function : \[f(x)=2\sqrt{x}\] square both sides and get \[\begin{align}(f(x))^2 &= (2\sqrt{x})^2\\~\\f(x)^2&=4x\end{align}\]

ganeshie8 (ganeshie8):

divide "4" both sides, what do u get ?

OpenStudy (anonymous):

2x

ganeshie8 (ganeshie8):

nope try again

OpenStudy (anonymous):

I'm sorry i can't seem to do it

ganeshie8 (ganeshie8):

thats okay, let me show u how to do it this one time We have : \[\begin{align}f(x)^2&=4x\end{align}\] Dividing "4" both sides : \[\begin{align}\frac{f(x)^2}{4}&=\frac{4x}{4}\end{align}\] On right hand side, we can cancel "4" top and bottom and get \[\begin{align}\frac{f(x)^2}{4}&=x\end{align}\]

ganeshie8 (ganeshie8):

which is same as \[x = \dfrac{f(x)^2}{4}\]

ganeshie8 (ganeshie8):

plug that in g(x) in place of "x" everywhere simplify

ganeshie8 (ganeshie8):

\[\begin{align} g(\color{blue}{x})&=\color{blue}{x}^2+\color{blue}{x}\\~\\ &=\left(\color{blue}{\frac{f(x)^2}{4}}\right)^2+\color{blue}{\frac{f(x)^2}{4}}\\~\\ &=\color{blue}{\frac{f(x)^4}{16}} + \color{blue}{\frac{f(x)^2}{4}}\\~\\ \end{align}\]

OpenStudy (anonymous):

so that's the answer right? @ganeshie8